tell me proof for this formula this formula in maths
(a+b)2=

Grade:12

14 Answers

N JYOTHEESWAR
8 years ago
a2+2ab+b2  
this is the expansion form of the formulae of the above question.so use the above formulae whenever you want to know or solve the problem
 
SAI SARDAR
8 years ago
(a+b)2=a2+b2+2ab
you devide (a+b)2 into (a+b)(a+b) and you can multiply then you will definetly got above one.
                          Thank you.
T.kumar
8 years ago
a+b)2=a2+b2+2ab
you devide (a+b)2 into (a+b)(a+b) and you can multiply then you will definetly got above one.
                          Thank you.
Prabhakar ch
8 years ago
If you multiply (a+b) and (a+b) then it will come. if multiply tha two it will come a2+b2+2ab.So please multiply tha two and answer will come
vannala shivanand
8 years ago
first draw a line with a point which divides a,b . then total distance of that line is a+b. the square of this disatnce is (a+b)2 .and it is equal to a2 +b2 +2ab.hence it is proved.
 
NARAYANA
8 years ago
WE REWRITE THE EQUATION (A+B)^2 IS (A+B)*(A+B) THEN WE GET ANSWER IS A^2+B^2+2AB THERE IS ANOTHER WAY TO SOLVE THIS BUT THIS IS SIMPLE WAY.
Dheeru chowdary
8 years ago
The expansion of the given algebraic formula is a^2 +b^2 + 2ab
   & by simple expansion u will get this ,,
RAMCHANDRARAO
8 years ago
(A+B)2 IS DIVIDED INTO (A+B)*(A+B)NOW YOU CAN MULTIPLY THESE TWO YOU DEFINETLY GOT    
(A+B)2 =A2+B2+2AB THANK YOU
NAGOORVALI
8 years ago
 
(a+b)2=a^2+b^2+2ab this is form in binomial expansion if interested you invent differnt modals we also know pascal triangle
Gowri sankar
8 years ago
DEAR CHETHAN,
                      (a+b)2=a2+b2+2ab this is form in binomial expression if interested you invent different modals we also known pascal triangle.
 
mohan
8 years ago
dear iitian first draw a line with a point which divides a,b . then total distance of that line is a+b. the square of this disatnce is (a+b)2 .and it is equal to a2 +b2 +2ab.hence it is proved.
Y MANJUNADH
8 years ago
dear chetan the proof of (a+b) 2=a2+b2+2ab it is a very usefull formula for solving problems.
and is called the binomial expansion formula
sreekanth sree
8 years ago
hello chethan (a+b)2=a2+B2 +2ab is the answer of this question.it is very important formula.(a+b)*(a+b)=a2+B2 +2ab….….….….….….….….….….….….….…...
manichandra
8 years ago
 
(a+b)2=a2+b2+2ab
you devide (a+b)2 into (a+b)(a+b) and you can multiply then you will definetly got above one.
                          Thank you.